Raul wants to join the directors of World Trade but currently owns no shares in the company. He knows that no one else will help elect him. Assume there are 46,000 shares outstanding at a market price of $12.80 a share. What is the minimum amount Raul must spend to acquire a seat on the board of directors if there are three open seats and straight voting applies?

A. $294,400.00

B. $196,266.67

C. $147,212.80

D. $294,412.80

E. $147,200.00
Answer #1

Answer:-

Raul must own more than 50 % of the total number of outstanding shares to acquire a seat on the board of directors if there are three open seats and straight voting applies.

Here the number of shares outstanding = 46000
Therefore total number of shares required = (46000 /2 ) + 1 = 23000 + 1 = 23001 shares
market price of share = $ 12.80

Total amount = 23001 x $ 12,80 = $ 294412.8 ( The amount that Raul needs to spend to acquire a seat on board of directors)

Therefore Option D is correct.